# 3DViT-GAT: A Unified Atlas-Based 3D Vision Transformer and Graph Learning Framework for Major Depressive Disorder Detection Using Structural MRI Data

### v1

@@ Abstract
Major depressive disorder (MDD) is a prevalent mental health condition that negatively impacts both individual well-being and global public health. Automated detection of MDD using structural magnetic resonance imaging (sMRI) and deep learning (DL) methods holds increasing promise for improving diagnostic accuracy and enabling early intervention. Most existing methods employ either voxel-level features or handcrafted regional representations built from predefined brain atlases, limiting their ability to capture complex brain patterns. This paper develops a unified pipeline that utilizes Vision Transformers (ViTs) for extracting 3D region embeddings from sMRI data and Graph Neural Network (GNN) for classification. We explore two strategies for defining regions: (1) an atlas-based approach using predefined structural and functional brain atlases, and (2) an cube-based method by which ViTs are trained directly to identify regions from uniformly extracted 3D patches. Further, cosine similarity graphs are generated to model interregional relationships, and guide GNN-based classification. Extensive experiments were conducted using the REST-meta-MDD dataset to demonstrate the effectiveness of our model. With stratified 10-fold cross-validation, the best model obtained 78.98% accuracy, 76.54% sensitivity, 81.58% specificity, 81.58% precision, and 78.98% F1-score. Further, atlas-based models consistently outperformed the cube-based approach, highlighting the importance of using domain-specific anatomical priors for MDD detection.
@@ 3.0.2 DATA PREPROCESSING
The T1-weighted sMRI scans were preprocessed using Diffeomorphic Anatomical Registration Through Exponentiated Lie algebra (DPARSF) software following the YAN processing protocol [20]. Preprocessing involved discarding the first ten volumes to ensure magnetization equilibrium, correcting head motion, segmenting into gray matter (GM), white matter (WM), and cerebrospinal fluid (CF), normalizing to the Montreal Neurological Institute (MNI) template, and spatial smoothing. Additionally, sites with fewer than ten subjects were excluded, and covariates were regressed to ensure data quality and consistency [20]. The resulting sMRI dataset contained 1586 subjects.
Notably, we performed additional preprocessing steps tailored to atlas-based analysis in our previous study using fMRI of the same cohort [21]. Specifically, we excluded subjects with incomplete or missing signal values in certain brain regions to ensure regional consistency across the dataset. This refinement resulted in 1563 subjects, of which 810 were MDD patients and 753 were HC. Our current study utilized the same refined subset of 1,563 subjects from our previous fMRI analysis of REST-meta-MDD. This permits fair cross-modal comparisons and enables the development of robust multimodal diagnostic frameworks that integrate structural and functional MRI modalities. Further, we utilized GM volumes in the MNI space to represent the original sMRI data. Each sMRI image had a standardized spatial dimension of voxels. A detailed description of the demographic characteristics of the subjects included in our study is provided in Table 1.
@@ 5.2 EXPERIMENTAL RESULTS
In this section, we conduct an extensive analysis to evaluate the effectiveness of our proposed 3DViT-GAT framework under several experimental configurations. Three main scenarios are considered in our investigation. First, we assess the standalone performance of the 3D ViT model using two distinct input paradigms: (1) atlas-based extraction, which guides region segmentation using brain atlases, and (2) cube-based extraction, which divides the brain into uniform, non-overlapping 3D cubes. Second, our full 3DViT-GAT hybrid model is constructed by integrating ViT-derived embeddings into a GAT. Graph-based classification is similarly investigated utilizing both atlas-based and cube-based graph representations. Finally, we compare the performance of our model with some of the state-of-the-art methods to demonstrate its competitiveness and generalizability for detecting MDD based on sMRI data.
@@ 5.2.1 RESULTS OF 3D VIT-BASED MODELS
Table 4 summarizes the average classification performance of the 3D ViT-based models across 10-fold cross-validation. Among the atlas-based models, the HO-based model achieved the highest accuracy of 73.442.15%, along with strong performance in sensitivity of 77.284.39%, specificity of 69.346.12%, precision of 73.133.43%, and F1-score of 74.991.95%. Meanwhile, the AAL-based model produced the highest F1-score of 75.541.25%, indicating a good balance between sensitivity and precision. However, it also exhibited the lowest specificity of 66.365.93% and a relatively lower precision of 71.852.85% compared to other atlas-based models. The Dose-based model scored highest in both specificity (69.616.35%) and precision (73.233.45%), indicating its strength in correctly identifying negative cases (HCs) and minimizing false positives. The CK-based model yielded 73.251.59% accuracy, 77.907.80% sensitivity, 68.298.12% specificity, 72.823.60% precision, and 74.882.51% F1-score. In addition, the comparison revealed that the cube-based model underperformed all atlas-based models in three metrics: accuracy of 72.291.89%, specificity of 63.167.94%, and precision of 70.362.88%.
In addition, the comparison revealed that the cube-based model underperformed relative to all atlas-based models in terms of accuracy, specificity, and precision. Specifically, it achieved an accuracy of 72.291.89%, a specificity of 63.167.94%, and a precision of 70.362.88%. These findings indicate that the model is unable to accurately identify true negatives and positive predictive instances. Despite these limitations, the cube-based model exhibited the highest sensitivity among all models, reaching 80.867.60%, reflecting its greater ability to detect true positive cases. However, this improvement came at the expense of increased false positives, as reflected in its lower specificity and precision. In contrast, the atlas-based models showed more stable and balanced performance across all evaluation metrics, making them more suitable for real-world diagnostic settings where balancing sensitivity and specificity is crucial. Overall, these findings provide evidence for the superior discriminatory power of atlas-based region extraction over cube-based partitioning, which uses anatomical brain partitioning to provide more informative and robust feature representations for ViT-based MDD classification.
Moreover, Table 5 displays the results obtained by the fold that achieved the highest accuracy. In the Ho-based model, the best model among all 10 folds produced 77.71% accuracy, 80.25% sensitivity, 75.00% specificity, 77.38% precision, and 78.79% F1-score. The best AAL-based model yielded 77.07% accuracy, 77.78% sensitivity, 76.74% specificity, 77.78% precision, and 77.78% F1-score. The best CK-based model had 76.43% accuracy, 83.95% sensitivity, 68.42% specificity, 73.91% precision, and 78.61% F1-score. Similarly, the best Dose-based model achieved 76.43% accuracy, 81.48% sensitivity, 71.05% specificity, 75.00% precision, and 78.11% F1-score. While the best cube-based model had 74.52% accuracy, 83.95% sensitivity, 64.47% specificity, 71.58% precision, and 77.27% F1-score.
@@ 5.2.2 RESULTS OF 3DVIT-GAT MODELS
Among all evaluated models, our HO-based 3DViT-GAT model achieved the best performance, with accuracy of 77.770.66%, sensitivity of 74.322.05%, specificity of 81.451.71%, precision of 81.061.12%, and F1-score of 77.520.91%. These results demonstrate solid and balanced classification capabilities across all evaluation metrics, emphasizing the advantages of combining atlas-guided region extraction with ViT-based representations and graph learning. The model proposed by Xiao et al. [18] yielded the second highest performance, with 77.17% accuracy, 80.30% sensitivity, 74.00% specificity, 75.56%, and 77.86% F1-score. Although this model showed a strong performance, their Transformer-based architecture operated on uniformly segmented 3D blocks without anatomical guidance, limiting its ability to capture fine-grained brain structures relevant to accurate diagnosis. Gao et al. [17] constructed structural covariance matrices using the AAL atlas and employed a 3D CNN to heatmap-enhanced GM images, attained accuracy of 76.64%, sensitivity of 73.52%, specificity of 79.40%, precision of 88.73%, and F1-score of 80.41%. Although it has a high F1-score and precision, its low sensitivity reflects its limited ability to detect true MDD cases, which can be costly in clinical settings.
@@
The Dose-based model also showed strong performance, with 75.991.18% accuracy, 70.395.80% specificity, 74.773.08% precision, and 77.730.69% F1-score. It also exhibited a sensitivity of 81.233.70%, which is higher than that of its competitors. Similarly, the Cube-based model reported 74.960.50% accuracy, 76.303.70% sensitivity, 73.553.88% specificity, 75.591.86% precision, and 75.841.02% F1-score. However, its lack of anatomical localization may limit its ability to detect meaningful morphological patterns critical for MDD detection. Notably, most atlas-guided models (e.g., AAL, HO, and Dose) performed better than the Cube-based model on almost all evaluation metrics, demonstrating the importance of anatomical guidance in improving representation learning. By contrast, the CK-based model provided the lowest performance among all atlas-guided strategies, with accuracy of 72.801.28%, sensitivity of 75.063.26%, specificity of 70.403.07%, precision of 73.051.64%, and F1-score of 73.981.53%.
@@
To support these findings, we performed two-sample t-tests (Table 7) comparing each atlas-based model with the Cube-based model. This study assumed that statistically significant differences existed between models when . The results showed that there was a significant difference in sensitivity, specificity, and precision between the AAL-based model and the Cube-based model. Meanwhile, there were significant differences between the HO-based model and the Cube-based model in terms of accuracy, specificity, precision, and F1-score. Moreover, the Dose-based model and Cube-based model showed significant differences in accuracy, sensitivity, and F1-score. There was also a significant difference in accuracy, precision, and F1-score between the Cube-based and CK-based models, supporting the superiority of the Cube-based approach over CK. Nevertheless, the Cube-based model performed significantly worse than all other atlas-based models, highlighting the significance of structured, anatomically guided ROI extraction to improve classification performance.
@@
Furthermore, Table 8 illustrates the evaluation metrics for the fold with the highest accuracy for the 3DViT-GAT models. The best performing Ho-based model produced 78.98% accuracy, 76.54% sensitivity, 81.58% specificity, 81.58% precision, and 78.98% F1-score. In Dose-based model, the best model achieved accuracy of 77.71%, sensitivity of 75.31%, specificity of 80.26%, precision of 80.26%, and F1-score of 77.71%. The best AAL-based model had 76.43% accuracy, 80.25% sensitivity, 72.37% specificity, 75.58% precision, and 77.84% F1-score. In addition, the best Cube-based model recorded 75.80% accuracy, 74.07% sensitivity, 77.64% specificity, 77.92% precision, and 75.95% F1-score. While the best CK-based model had 74.52% accuracy, 79.01% sensitivity, 69.74% specificity, 73.56% precision, and 76.19% F1-score.

### v3

@@ Abstract
Major depressive disorder (MDD) is a prevalent mental health condition that negatively impacts both individual well-being and global public health. Automated detection of MDD using structural magnetic resonance imaging (sMRI) and deep learning (DL) methods holds increasing promise for improving diagnostic accuracy and enabling early intervention. Most existing methods employ either voxel-level features or handcrafted regional representations built from predefined brain atlases, limiting their ability to capture complex brain patterns. This paper develops a unified pipeline that utilizes Vision Transformers (ViTs) for extracting 3D region embeddings from sMRI data and Graph Neural Network (GNN) for classification. We explore two strategies for defining regions: (1) an atlas-based approach using predefined structural and functional brain atlases, and (2) an cube-based method by which ViTs are trained directly to identify regions from uniformly extracted 3D patches. Further, cosine similarity graphs are generated to model interregional relationships, and guide GNN-based classification. Extensive experiments were conducted using the REST-meta-MDD dataset to demonstrate the effectiveness of our model. With stratified 10-fold cross-validation, the best model obtained 81.51% accuracy, 85.94% sensitivity, 76.36% specificity, 80.88% precision, and 83.33% F1-score. Further, atlas-based models consistently outperformed the cube-based approach, highlighting the importance of using domain-specific anatomical priors for MDD detection.
@@ 3.0.2 DATA PREPROCESSING
The T1-weighted sMRI scans were preprocessed using Diffeomorphic Anatomical Registration Through Exponentiated Lie algebra (DPARSF) software [20]. Preprocessing involved discarding the first ten volumes to ensure magnetization equilibrium, correcting head motion, segmenting into gray matter (GM), white matter (WM), and cerebrospinal fluid (CF), normalizing to the Montreal Neurological Institute (MNI) template, and spatial smoothing. Additionally, a duplicate site was discovered after consortium data sharing and discarded, and covariates were regressed to ensure data quality and consistency [20]. The resulting sMRI dataset contained 1586 subjects, of which 1276 were MDD patients and 1104 were HC.
Further, we utilized GM volumes in the MNI space to represent the original sMRI data. Each sMRI image had a standardized spatial dimension of voxels. A detailed description of the demographic characteristics of the subjects included in our study is provided in Table 1. .
@@ 5.2 EXPERIMENTAL RESULTS
In this section, we conduct an extensive analysis to evaluate the effectiveness of our proposed 3DViT-GAT framework under several experimental configurations. Three main scenarios are considered in our investigation. First, we assess the standalone performance of the 3D ViT model using two distinct input paradigms: (1) atlas-based extraction, which guides region segmentation using brain atlases, and (2) cube-based extraction, which divides the brain into uniform, non-overlapping 3D cubes. Second, our full 3DViT-GAT hybrid model is constructed by integrating ViT-derived embeddings into a GAT. Graph-based classification is similarly investigated utilizing both atlas-based and cube-based graph representations. Third, we compare the performance of our model with some of the state-of-the-art methods to demonstrate its competitiveness and generalizability for detecting MDD based on sMRI data. Finally, a model explainability analysis are conducted to identify the most critical brain ROIs associated with MDD, HC, and MDD vs. HC.
@@ 5.2.1 RESULTS OF 3D VIT-BASED MODELS
Table 4 summarizes the average classification performance of the 3D ViT-based models across 10-fold cross-validation. Among the tested models, the HO-based model achieved the highest accuracy of 76.930.83%, along with the highest specificity (73.366.58%) and precision (78.093.34%), indicating its strength in correctly identifying negative cases (HC) and minimizing false positives. In addition, it achieved strong performance in terms of sensitivity (80.005.07%) and F1-score (78.821.11%). The Cube-based model produced the second highest accuracy at 76.811.27% despite the absence of anatomical priors. It also exhibited the highest sensitivity (83.676.28%) and F1-score (79.451.39%), reflecting its greater ability to detect true positive cases (MDD). However, this improvement came at the expense of increased false positives, as reflected in its lower specificity (68.828.26%) and precision (76.163.59%). The Dose-based model yielded 76.721.45% accuracy, 82.193.69% sensitivity, 70.366.14% specificity, 76.583.08% precision, and 79.151.05% F1-score. Meanwhile, the AAL-based model attained 76.600.98% accuracy, 81.565.80% sensitivity, 70.826.84% specificity, 76.803.07% precision, and 78.881.42% F1-score. Furthermore, The CK-based model delivered comparable performance, with an accuracy of 76.391.40%, a sensitivity of 80.553.23%, a specificity of 71.555.93%, a precision of 76.953.17%, and an F1-score of 78.590.87%.
Moreover, Table 5 displays the results obtained by the fold that achieved the highest accuracy. Among the atlas-based models,the best Dose-based model among all 10 folds achieved 78.99% accuracy, 78.91% sensitivity, 79.09% specificity, 81.45% precision, and 80.16% F1-score. The best CK-based model had 78.15% accuracy, 81.25% sensitivity, 74.55% specificity, 78.79% precision, and 80.00% F1-score. In the AAL-based model, the best model produced 77.73% accuracy, 92.97% sensitivity, 60.00% specificity, 73.01% precision, and 81.79% F1-score. Similarly, the best HO-based model yielded 77.73% accuracy, 83.59% sensitivity, 70.91% specificity, 76.98% precision, and 80.15% F1-score. On the other hand, the best cube-based model had 78.57% accuracy, 87.50% sensitivity, 68.18% specificity, 76.19% precision, and 81.45% F1-score.
In summary, the atlas-based models showed more stable and balanced performance across all evaluation metrics, making them more suitable for real-world diagnostic settings where balancing sensitivity and specificity is crucial. Overall, these findings provide evidence for the superior discriminatory power of atlas-based region extraction over cube-based partitioning, which uses anatomical brain partitioning to provide more informative and robust feature representations for ViT-based MDD classification.
@@ 5.2.2 RESULTS OF 3DVIT-GAT MODELS
Among all evaluated models, our Dose-based 3DViT-GAT model achieved the best performance, with an accuracy of 80.880.39%, a sensitivity of 86.252.91%, a specificity of 74.643.48%, a precision of 79.921.69%, and an F1-score of 82.900.57%. The AAl-based achieved the second highest performance, with an accuracy of 79.710.42%, a specificity of 70.552.31%, a precision of 77.611.04%, and an F1-score of 82.270.42%. It also exhibited a sensitivity of 87.581.80%, which is higher than that of its competitors. Similarly, the CK-based model produced 79.200.60% accuracy, 83.671.86% sensitivity, 74.002.08% specificity, 78.951.05% precision, and 81.220.64% F1-score. These results demonstrate that our proposed atlas-based models provide solid and balanced classification capabilities across all evaluation metrics, emphasizing the advantages of combining atlas-guided region extraction with ViT-based representations and graph learning. In contrast, the Cube-based model reported 78.951.25% accuracy, 82.032.99% sensitivity, 75.363.87% specificity, 79.592.19% precision, and 80.731.20% F1-score. However, its lack of anatomical localization may limit its ability to detect meaningful morphological patterns critical for MDD detection.
The model proposed by Xiao et al. [18] yielded 77.17% accuracy, 80.30% sensitivity, 74.00% specificity, 75.56% precision, and 77.86% F1-score. Although this model showed a strong performance, their Transformer-based architecture operated on uniformly segmented 3D blocks without anatomical guidance, limiting its ability to capture fine-grained brain structures relevant to accurate diagnosis. Our HO-based model also showed strong performance, with an accuracy of 77.060.80%, a sensitivity of 76.095.45%, a specificity of 78.187.14%, a precision of 80.693.89%, and an F1-score of 78.051.14%. Gao et al. [17] constructed structural covariance matrices using the AAL atlas and employed a 3D CNN to heatmap-enhanced GM images, attained an accuracy of 76.64%, a sensitivity of 73.52%, a specificity of 79.40%, a precision of 88.73%, and an F1-score of 80.41%. Although it has high specificity and precision, its low sensitivity reflects its limited ability to detect true MDD cases, which can be costly in clinical settings.
@@
To support these findings, we performed two-sample t-tests (Table 7) comparing each atlas-based model with the Cube-based model. This study assumed that statistically significant differences existed between models when . The results showed that there was a significant difference in sensitivity, specificity, precision, and F1-score between the AAL-based model and the Cube-based model. Meanwhile, there were significant differences between the Dose-based model and the Cube-based model in terms of in accuracy, sensitivity, and F1-score. There was also a significant difference in accuracy, sensitivity, and F1-score between the Cube-based and HO-based models, supporting the superiority of the Cube-based approach over HO. Nevertheless, the Cube-based model performed significantly worse than all other atlas-based models, highlighting the significance of structured, anatomically guided ROI extraction to improve classification performance. Additionally, there is no significant difference between the Cube-based and CK-based models.
@@
Furthermore, Table 8 illustrates the evaluation metrics for the fold with the highest accuracy for the 3DViT-GAT models. The best performing Dose-based model produced 81.51% accuracy, 85.94% sensitivity, 76.36% specificity, 80.88% precision, and 83.33% F1-score. In AAL-based model, the best model achieved an accuracy of 80.25%, a sensitivity of 90.63%, a specificity of 68.18%, a precision of 76.82%, and an F1-score of 83.15%. The best CK-based model had 80.25% accuracy, 87.50% sensitivity, 71.82% specificity, 78.32% precision, and 82.66% F1-score. In addition, the best HO-based model recorded 78.15% accuracy, 78.13% sensitivity, 78.18% specificity, 80.65% precision, and 79.37% F1-score. While, the best Cube-based model had 81.51% accuracy, 81.25% sensitivity, 81.82% specificity, 83.87% precision, and 82.54% F1-score.
@@
5.2.3 Model Explainability And Interpretation
To address our concerns about the interpretability and biological relevance of the extracted brain regions, we performed a post-hoc explainability analysis to determine which brain regions played the most significant role in the classification of MDD and HC. To achieve this, we applied a node-level interpretability approach derived from GNNExplainer [30]. Moreover, the explainability analysis are conducted using the two best-performing atlas-based 3DViT-GAT models (Dose and AAL) in our study to ensure meaningful and representative results. For each trained model, the contribution of each node (ROI) is calculated across all subjects in the test set, separately for MDD and HC classes, as well as for the overall classification. The ROI importance scores are then averaged over all test subjects to derive reliable estimates of regional significance. We visualized the top ten ROIs on a cortical surface template using bubble plots, with bubble size and color intensity representing the importance of each ROI.
Figure 2 (a-c) provides the explainability results derived from the Dose-based model, highlighting the most salient ROIs contributing to the classification of MDD, HC, and MDD vs. HC, respectively. All three explainability maps indicated that highly influential ROIs are located within the fronto-parietal network (FPN), occipital network (ON), cerebellum network (CN), sensorimotor network (SMN), and cingulo-opercular network (CON). A summary of the top ten salient ROIs identified within these networks across all three analyses can be found in Table 9.
Subsequently, the AAL-based model is examined to identify regions that contributed to the prediction outcomes. Notably, the AAL atlas differs from the Dose atlas in that brain regions are grouped anatomically into distinct lobes rather than functional networks [21]. Figure 3 presents explainability results for the AAL-based model, emphasizing the most significant ROIs contributing to MDD, HC, and MDD vs. HC classifications. Using this model, the explainability maps revealed that the most salient ROIs are located in the frontal, insula, temporal, parietal, and occipital lobes of the brain. Table 10 summarizes the ten most significant ROIs within these lobes.
| Rank | MDD | HC | MDD vs. HC | |||
| Brain Region | Network | Brain Region | Network | Brain Region | Network | |
| 1 | R dlPFC | FPN | R IPL | FPN | R IPL | FPN |
| 2 | R IPL | FPN | L post occipital | ON | R dlPFC | FPN |
| 3 | L inf cerebellum | CN | L inf cerebellum | CN | L post occipital | ON |
| 4 | L post occipital | ON | L precentral gyrus | SMN | L inf cerebellum | CN |
| 5 | L precentral gyrus | SMN | R dlPFC | FPN | L precentral gyrus | SMN |
| 6 | L basal ganglia | CON | L precentral gyrus | SMN | L precentral gyrus | SMN |
| 7 | L vFC | SMN | L vFC | SMN | L vFC | SMN |
| 8 | L precentral gyrus | SMN | L basal ganglia | CON | L basal ganglia | CON |
| 9 | L inf cerebellum | CN | L inf cerebellum | CN | L inf cerebellum | CN |
| 10 | R parietal | SMN | R dFC | FPN | R parietal | SMN |
Abbreviations: MDD, major depressive disorder; HC, healthy control; L, left hemisphere; R, right hemisphere; dlPFC, dorsolateral prefrontal cortex; FPN, fronto-parietal network; IPL, inferior parietal lobule; inf, inferior; CN, cerebellum network; ON, occipital network; SMN, sensorimotor network; CON, cingulo-opercular network; vFC, ventral frontal cortex; dFC, dorsolateral frontal cortex.
| Rank | MDD | HC | MDD vs. HC | |||
| Brain Region | Lobe | Brain Region | Lobe | Brain Region | Lobe | |
| 1 | L PreCG | Frontal | R insula | Insula | R insula | Insula |
| 2 | R insula | Insula | R ITG | Temporal | L rectus | Frontal |
| 3 | L TPOmid | Temporal | R ORBinf | Frontal | L MFG | Frontal |
| 4 | R SPG | Parietal | L rectus | Frontal | R SPG | Parietal |
| 5 | L rectus | Frontal | L MFG | Frontal | R ITG | Temporal |
| 6 | L MFG | Frontal | R MOG | Occipital | R MOG | Occipital |
| 7 | R MOG | Occipital | R SPG | Parietal | L PreCG | Frontal |
| 8 | R olfactory | Frontal | R TPOsup | Temporal | L heschl | Temporal |
| 9 | L TPOsup | Temporal | L heschl | Temporal | L TPOmid | Temporal |
| 10 | L heschl | Temporal | R calcarine | Occipital | R olfactory | Frontal |
Abbreviations: MDD, major depressive disorder; HC, healthy control; L, left hemisphere; R, right hemisphere; PreCG, precentral gyrus; TPOmid, Temporal Middle Pole; SPG, Superior parietal gyrus; MFG, Middle frontal gyrus; MOG, Middle occipital gyrus; TPOsup, Temporal superior pole; ITG, inferior temporal gyrus; ORBinf, Orbitofrontal inferior cortex.